tom is 7 years older than maria. In five years, tom willl be one less twice maris's age. how old is maria
Found 2 solutions by ankor@dixie-net.com, josgarithmetic:
Answer by ankor@dixie-net.com(22740)

Tom is 7 years older than Maria.
t = m + 7
In five years, Tom will be one less twice Maris's age.
t + 5 = 2(m+5) - 1
t + 5 = 2m + 10 - 1
t = 2m + 9 - 5
t = 2m + 4
how old is Maria
Replace T with (m+7)
m + 7 = 2m + 4
7 - 4 = 2m - m
m = 3 yrs is Maria's age
:
:
Check this in the statement (Tom is 10)
"In five years, tom will be one less twice maris's age.
10 + 5 = 2(3+5) - 1)
15 = 2(8) - 1

Answer by josgarithmetic(39616)